Dolor had a sister who is 2 years older than twice her age the sum of their ages is 23 how old is spike
Answer by ankor@dixie-net.com(22740)

Dolor had a sister who is 2 years older than twice her age
2d = s - 2
2d + 2 = s
the sum of their ages is 23
d + s = 23
replace s with (2d+2)
d + 2d + 2 = 23
d + 2d = 23 - 2
3d = 21
d = 21/3
d = 7 years is Dolor's age
and
s = 2(7) + 2
s = 16 yrs is sister's age
:
Spike's age? I am not sure.